OpenStudy (anonymous):

Simplify: 5 + 2 radical 7 quantity squared

OpenStudy (lgbasallote):

\(\large (5 + \sqrt {7}) ^{2}\) is this the question?

OpenStudy (lgbasallote):

there's a 2 before 7...

OpenStudy (anonymous):

yes there is

OpenStudy (lgbasallote):

so \(\large (5 + 2\sqrt {7})^{2}\) is the problem right??

OpenStudy (anonymous):

yes it is

OpenStudy (lgbasallote):

you can use good ol FOIL \( (5 + 2\sqrt 7)(5+ 2\sqrt 7) = (5)(5) + 2(5)(2\sqrt 7) + (2 \sqrt 7)^2\) can you try solving that now?

OpenStudy (anonymous):

no lol

OpenStudy (lgbasallote):

what is 5 x 5?